# Exploring the Robustness of Language Models for Tabular Question Answering via Attention Analysis

### v3

@@ H.2 Correlation within individual attention
Figure 17: Heatmap showing Spearman correlation between changes in attention entropy and EM difference across all attention heads and layers in the LLama-8B-Instruct model on the WTQ dataset, unsorted to depict actual correlation over each attention head.
Figure 18: Scatter plot showing the correlation between changes in attention entropy and EM scores across five structural perturbations for the WTQ dataset. The plot highlights a strong positive correlation for Layer 0, Head 10, indicating high sensitivity of this head to structural perturbations from 17
Figure 19: Scatter plot showing the correlation between changes in attention entropy and EM scores across five structural perturbations for the WTQ dataset. In contrast to the previous head, Layer 19, Head 23 shows minimal correlation from 17
We find that structural perturbations in tabular inputs affect the internal attention dynamics of the LLMs specifically, Llama-8B-Instruct model on WTQ, and how those dynamics relate to the performance. Figure 17 presents an unsorted Spearman correlation heatmap between per-head changes in attention entropy and the corresponding EM score differences, revealing substantial heterogeneity across all 32 heads and 32 layers. To unpack this variability, Figures 18 and 19 zoom in on the two extremal heads identified by the heatmap: Layer 00, Head $10$ exhibits a strong positive relationship, higher entropy shifts reliably predict larger drops in EM, whereas Layer 19, Head 23 shows significantly less correlation, indicating that its attention distribution remains largely impervious to structural pertubation. Together, these visualizations underscore that certain attention heads are sensitive for model robustness, while others maintain stable focus patterns even under severe perturbations.

### v4

@@ H.1 Spearman Correlation within All Attention Heads
Figure 14: Heatmap showing Spearman correlation between changes in attention entropy and EM difference across all attention heads and layers in the Llama2-7B model(a, b and c) and Llama2-7B-chat model(d, e and f).
Figure 15: Heatmap showing Spearman correlation between changes in attention entropy and EM difference across all attention heads and layers in the Llama3-8B model(a, b and c) and Llama3-8B-instruct model(d, e and f).
Figure 16: Heatmap showing Spearman correlation between changes in attention entropy and EM difference across all attention heads and layers in the Mistral-7B model(a, b and c) and Mistral-7B-instruct model(d, e and f).
Figure 17: Heatmap showing Spearman correlation between changes in attention entropy and EM difference across all attention heads and layers in the Qwen2.5-7B model(a, b and c) and Qwen2.5-7B-instruct model(d, e and f).
Figure 18: Heatmap showing Spearman correlation between changes in attention entropy and EM difference across all attention heads and layers in the Qwen3-8B base model(a, b and c) and Qwen3-8B model(d, e and f).
Across both the model families Llama3 and Mistral, the middle layers consistently exhibit the strongest positive Spearman correlations between perturbation-induced changes in attention entropy and EM degradation. This recurring ‘hot spot’ in the mid-layer shows a general architectural property; middle transformer blocks appear to serve as critical junctions where structural perturbations most directly translate into downstream performance loss. Such robustness vulnerabilities likely stem from these layers’ dual role in integrating lower-level token interactions and preparing higher-level semantic abstractions, making them both information-rich and sensitive to distributional shifts.
However, differences emerge when contrasting base versus chat- or instruction-tuned variants. In the base models (Llama3-8B, Mistral-7B; Figure ( 15, and 16) subfigures a-c), the correlation patterns outside the middle layers fluctuate markedly, with a mixture of weak or even negative correlation, especially apparent on TAT-QA and SCITAB tasks. Specifically, for Figure 16(c),Mistral-7B does not correlate because for all the questions EM was 0. These oscillations align with the base models’ generally lower EM scores on these datasets, suggesting that noisy or unreliable predictions can obscure the coherent relationships between entropy and performance.
By contrast, the chat and instruct-adapted models (Llama3-8B-Instruct, Mistral-7B-Instruct; Figure (15 and 16) subfigures d-f) display stronger positive correlation, with some extending beyond the middle layers. On WTQ in particular, both the initial encoding layers (layers 020-2) and the highest layers (uppermost 293129-31 blocks) contribute positively to the entropy–EM relation, suggesting that conversational and instruction tuning bolsters the model’s resilience to perturbations at both the token-embedding stage and the final consolidation stage. This extended positive band likely reflects enhanced parameter alignment across the architecture, enabling more stable information propagation even under input disruptions.
In general, these results yield two primary implications. First, conversational and instruction-tuning systematically extends the alignment between attention entropy instabilities and performance degradation across a broader portion of the transformer hierarchy, thereby allowing perturbation robustness not only in the mid layers but also at its boundaries. Second, task domain complexity governs which layers most critically underpin model stability: general-domain benchmarks (e.g., WTQ) draw on a broad spectrum of transformer depths, whereas specialized datasets remain predominantly reliant on central layers.
Across the Qwen model families, the correlation heatmaps in Figures 17 and 18 reveal patterns broadly consistent with those reported for Llama3 and Mistral in the main text. For both Qwen 2.5 and Qwen 3, the middle transformer layers consistently exhibit the strongest positive Spearman correlations between perturbation-induced changes in attention entropy and EM degradation, reinforcing the interpretation of these layers as structural “choke points” where disruptions to schema alignment most directly translate into performance loss. Base variants (Figures 17a–c, 18a–c) show more fluctuation outside the mid-layer band, including weak or negative correlations on TAT-QA and SCITAB, mirroring the instability seen in Llama3-8B and Mistral-7B. This instability often coincides with lower overall EM, as in Qwen 2.5-7B on SCITAB. In contrast, the instruction-tuned counterparts (Figures 17d–f, 18d–f) display stronger and more spatially extended positive correlations, with WTQ in particular showing elevated values not only in the middle layers but also in the initial embedding layers (0–2) and uppermost layers, suggesting improved perturbation robustness at both early encoding and final integration stages. These trends indicate that, for Qwen as for Llama3 and Mistral, instruction tuning enhances the architectural alignment of attention–performance relationships, while the dataset domain continues to determine how localized or distributed the layerwise sensitivity appears.